# On strictly elliptic K3 surfaces and del Pezzo surfaces

### v1

@@ 1. Introduction
Finally, it is noteworthy that, as a consequence of recent work [CM23] by Clingher and Malmendier, the considered elliptic fibrations are not jacobian, i.e., they do not admit sections. However, it is not difficult to observe (see §5) that they admit bisections which can be induced by the $(-1)$-curves in the associated del Pezzo surface (see Example 3.10). Despite the absence of jacobian elliptic fibrations, and consequently the inability to consider Weierstrass models, these K3 surfaces are quite special due to the fact that by [AN06, §2.8] they have finite automorphism groups. Remarkably, they fall within the recent classification by Roulleau [Rou22], where some explicit projective models are studied and where the full lattice $\operatorname{NS}(X)$ is described (see §4).
@@ Acknowledgements
The authors would like to sincerely thank Michela Artebani, Giacomo Mezzedemi, Cecília Salgado, and Giancarlo Urzúa for various fruitful discussions regarding the constructions used in this article. P.C. has been partially funded by Universidad de La Frontera, Proyecto DIM23-0001. P.M. has been partially funded by Fondecyt ANID Project 1231214. S.T. was partially supported by Fondecyt ANID Project 3210518. The authors express their gratitude to Ceciília Salgado for her generous hospitality extended to S.T. during his visit to the Bernoulli Institute at the University of Groningen.
@@ 2.1. Conic bundles and smooth del Pezzo surfaces
$D=4L-2\sum_{j\in J}E_{j}-\sum_{k\in K}E_{k}$, with $J\subset\{1,\ldots,7\},|J|=4$ and $K=\{1,\ldots,7\}\setminus J$
$D=5L-E_{i}-2\sum_{j\in J}E_{j}$, with $i\in\{1,\dots,7\}$ and $J=\{1,\ldots,7\}\setminus\{i\}$
@@
$D=4L-2\sum_{j\in J}E_{j}-\sum_{k\in K}E_{k}$, with $J,K\subset\{1,\ldots,8\},|J|=4,|K|=3,J\cap K=\emptyset$
$D=5L-3E_{i}-2\sum_{j\in J}E_{j}-\sum_{k\in K}E_{i_{k}}$, with $i\in\{1,\ldots,8\}$,
@@
with $J,K,R\subset\{1,\dots,8\},|J|=2,|K|=4,|R|=2,J\cap K=J\cap R=K\cap R=\emptyset$
@@
For each $d$, one gets the possibilities for $D$ of Table 1 solving the Diophantine equation system.
A straightforward combinatorial computation gives the total number of conic bundles $N$ for each $d$ depending on $n$. ∎
@@ 2.2. Non-symplectic involutions on K3 surfaces
Let $X$ be a projective K3 surface and $\varphi:X\to X$ be an automorphism of finite order $n$. Let us recall that $\varphi$ is symplectic (resp. purely non-symplectic) if $\varphi^{\ast}\omega=\omega$ (resp. $\varphi^{\ast}\omega=\lambda\omega$ for some $\lambda$ a primitive $n$-rooth of the unity), where $\omega$ is a non-degenerate holomorphic $2$-form on $X$ (i.e., $\operatorname{H}^{0}(X,\Omega_{X}^{2})=\mathbf{C}\cdot\omega$). Given a K3 surface $X$ and a non-symplectic involution $\iota$ on $X$, we denote $X^{\iota}$ the fixed locus of the involution and $*\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$ the invariant lattice, i.e.
@@
where $C_{g}$ is a smooth curve of genus $g$ and the $R_{i}$’s are rational curves. Moreover, $g=\frac{22-r-a}{2}$ and $k=\frac{r-a}{2}$, where $*r=\operatorname{\operatorname{rk}}\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$ and $a$ is the length of $*\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$, i.e., $2^{a}=|\operatorname{discr}\operatorname{\operatorname{NS}}(X)^{\iota^{*}}|$.
@@
As observed in [Nik83], the invariants $(r,a)$ allow to recover the topological invariants $(g,k)$ of the fixed locus $X^{\iota}$. Viceversa, if the pair $(g,k)$ is known, a third invariant $\delta$ is needed to identify uniquely $*\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$: $\delta=0$ if and only if $X^{\iota}\sim 0\text{ mod }2$ in $H_{2}(X,\mathbf{Z})$, otherwise $\delta=1$. The picture of all possible invariant $(r,a,\delta)$ is presented in [AN89, Figure 1]. For completeness, we include it in Figure 1.
{tikzpicture}
@@ 3. Strictly elliptic involutions
Finally, observe that by [CM23, Corollary 3.3], the case $(r,a,\delta)=(6,4,0)$ does not admit jacobian elliptic fibrations and therefore it is not considered in the analysis of [GS20] despite the presence of a rational curve in $X^{\iota}$. On the other hand, it follows from [Nak07, Table 9] that the quotient surface $Z=X/\langle\iota\rangle$ is a singular del Pezzo surface of Gorenstein index two that can be realized via a Sarkisov link starting from $\mathbf{P}(1,1,4)$ or as a hypersurface of degree $5$ in $\mathbf{P}(1,1,1,4)$ (see [Nak07, Proposition 7.4, 7.11]).
@@
Let $(X,\iota)$ be a pair of strictly elliptic type. Suppose that $C_{g}$ is the smooth irreducible curve of genus $2\leq g\leq 10$ fixed by $\iota$.
@@
First, let us consider $\mathcal{E}:X\to\mathbf{P}^{1}$ an arbitrary elliptic fibration. Since $(X,\iota)$ is strictly elliptic, we know that $\iota$ must map each fiber of $\mathcal{E}$ to itself (see Remark 3.3), and thus $\mathcal{E}$ factors through a fibration $f:X/\langle\iota\rangle\simeq Z_{d}\to\mathbf{P}^{1}$. Since $Z_{d}$ is a smooth del Pezzo surface and $f$ has connected fibers, it follows that $f$ is a $K_{Z_{d}}$-negative contraction with one-dimensional fibers and thus is a conic bundle by [And85, Theorem 3.1 (ii)].
@@
{tikzcd}
@@ 4. Examples and Néron-Severi lattices
$w^{2}=(a^{4}-a^{2}b^{2}+b^{4})s^{4}+\tfrac{1}{2}(-a^{3}b+a^{2}b^{2})s^{3}+\tfrac{1}{2}(a^{3}b+3a^{2}b^{2}+ab^{3}-4b^{4})s^{2}+({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}-}a^{2}b^{2}+b^{4}).$

### v2

@@ 1. Introduction
Finally, it is noteworthy that, as a consequence of recent work [CM23] by Clingher and Malmendier, the considered elliptic fibrations are not jacobian, i.e., they do not admit sections. However, it is not difficult to observe (see §5) that they admit bisections which can be induced by the $(-1)$-curves in the associated del Pezzo surface (see Example 3.10). Despite the absence of jacobian elliptic fibrations, and consequently the inability to consider Weierstrass models, these K3 surfaces are quite special due to the fact that by [AN06, §2.8] they have finite automorphism groups. Remarkably, they fall within the recent work by Roulleau [Rou22], where some explicit projective models are studied and where the full lattice $\operatorname{NS}(X)$ is described (see §4).
@@ Acknowledgements
The authors would like to sincerely thank Michela Artebani, Giacomo Mezzedemi, Cecília Salgado, and Giancarlo Urzúa for various fruitful discussions regarding the constructions used in this article. Special thanks go to Cinzia Casagrande, who informed us about reference [Der06] and allowed us to correct errors in the counting of conic classes at low degree. P.C. has been partially funded by Universidad de La Frontera, Proyecto DIM23-0001 and Fondecyt ANID Project 1240360. P.M. has been partially funded by Fondecyt ANID Projects 1231214 and 1240101. S.T. was partially supported by Fondecyt ANID Project 3210518. The authors express their gratitude to Cecília Salgado for her generous hospitality extended to S.T. during his visit to the Bernoulli Institute at the University of Groningen.
@@ 2.1. Conic bundles and smooth del Pezzo surfaces
$D=4L-2\sum_{j\in J}E_{j}-\sum_{k\in K}E_{k}$, $J\subset\{1,\ldots,7\},|J|=4$, $K=\{1,\ldots,7\}\setminus J$
$D=5L-E_{i}-2\sum_{j\in J}E_{j}$, $i\in\{1,\dots,7\}$, $J=\{1,\ldots,7\}\setminus\{i\}$
@@
$D=4L-2\sum_{j\in J}E_{j}-\sum_{k\in K}E_{k}$, $J,K\subset\{1,\ldots,8\},|J|=4,|K|=3,J\cap K=\emptyset$
$D=4L-3E_{i}-\sum_{j\in J}E_{j},\;i\in\{1,\ldots,8\},\;J=\{1,\ldots,8\}\setminus\{i\}$
$D=5L-3E_{i}-2\sum_{j\in J}E_{j}-\sum_{k\in K}E_{i_{k}}$, $i\in\{1,\ldots,8\}$,
@@
$J,K,R\subset\{1,\dots,8\},|J|=2,|K|=4,|R|=2,J\cap K=J\cap R=K\cap R=\emptyset$
$D=7L-4E_{i}-3E_{j}-2\sum_{k\in K}E_{k},\;i\in\{1,\ldots,8\},\;j\in\{1,\ldots,8\}\setminus\{i\},\;K=\{1,\ldots,8\}\setminus\{i,j\}$
$D=7L-E_{i}-2\sum_{j\in J}E_{j}-3\sum_{k\in K}E_{k},\;i\in\{1,\ldots,8\},\;J,K\subset\{1,\ldots,8\}\setminus\{i\},\;|J|=3,\;|K|=4,\;J\cap K=\emptyset$
$D=8L-4E_{i}-2\sum_{j\in J}E_{j}-3\sum_{k\in K}E_{k},\;i\in\{1,\ldots,8\},\;J,K\subset\{1,\ldots,8\}\setminus\{i\},\;|J|=3,\;|K|=4,\;J\cap K=\emptyset$
$D=8L-E_{i}-3\sum_{j\in J}E_{j},\;i\in\{1,\ldots,8\},\;J=\{1,\ldots,8\}\setminus\{i\}$
$D=9L-2E_{i}-4\sum_{j\in J}E_{j}-3\sum_{k\in K}E_{k},\;i\in\{1,\ldots,8\},\;J,K\subset\{1,\ldots,8\}\setminus\{i\},\;|J|=2,\;|K|=5,\;J\cap K=\emptyset$
$D=10L-3\sum_{i\in I}E_{i}-4\sum_{j\in J},\;I,J\subset\{1,\ldots,8\},\;|J|=|K|=4,\;J\cap K=\emptyset$
$D=11L-3E_{i}-4\sum_{j\in J}E_{j},\;i\in\{1,\ldots,8\},\;J=\{1,\ldots,8\}\setminus\{i\}$
@@
This is a classical fact that can be found in [Der06, §2] (see also [TVAV09, Table 2]). For the reader’s convenience, we give a self-contained proof.
@@
For each $d$, one gets the possibilities for $D$ of Table 1 solving the Diophantine equation system. A straightforward combinatorial computation gives the total number of conic bundles $N$ for each $d$ depending on $n$. ∎
@@ 2.2. Non-symplectic involutions on K3 surfaces
Let $X$ be a projective K3 surface and $\varphi:X\to X$ be an automorphism of finite order $n$. Let us recall that $\varphi$ is symplectic (resp. purely non-symplectic) if $\varphi^{\ast}\omega=\omega$ (resp. $\varphi^{\ast}\omega=\lambda\omega$ for some $\lambda$ a primitive $n$-rooth of the unity), where $\omega$ is a non-degenerate holomorphic $2$-form on $X$ (i.e., $\operatorname{H}^{0}(X,\Omega_{X}^{2})=\mathbf{C}\cdot\omega$). Given a K3 surface $X$ and a non-symplectic involution $\iota$ on $X$, we denote $X^{\iota}$ the fixed locus of the involution and $\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$ the invariant lattice, i.e.
@@
where $C_{g}$ is a smooth curve of genus $g$ and the $R_{i}$’s are rational curves. Moreover, $g=\frac{22-r-a}{2}$ and $k=\frac{r-a}{2}$, where $r=\operatorname{\operatorname{rk}}\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$ and $a$ is the length of $\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$, i.e., $2^{a}=|\operatorname{discr}\operatorname{\operatorname{NS}}(X)^{\iota^{*}}|$.
@@
As observed in [Nik83], the invariants $(r,a)$ allow to recover the topological invariants $(g,k)$ of the fixed locus $X^{\iota}$. Viceversa, if the pair $(g,k)$ is known, a third invariant $\delta$ is needed to identify uniquely $\operatorname{\operatorname{NS}}(X)^{\iota^{*}}$: $\delta=0$ if and only if $X^{\iota}\sim 0\text{ mod }2$ in $H_{2}(X,\mathbf{Z})$, otherwise $\delta=1$. The picture of all possible invariant $(r,a,\delta)$ is presented in [AN89, Figure 1]. For completeness, we include it in Figure 1.
[…]
@@ 3. Strictly elliptic involutions
Finally, observe that by [CM23, Corollary 3.3], the case $(r,a,\delta)=(6,4,0)$ does not admit jacobian elliptic fibrations and therefore it is not considered in the analysis of [GS20] despite the presence of a rational curve in $X^{\iota}$. On the other hand, by [Nak07, Table 9], $Z=\operatorname{Proj}\oplus_{m\geq 0}\operatorname{H}^{0}(X,\mathcal{O}_{X}(mC_{g}))^{\iota}$ is a singular del Pezzo surface of Gorenstein index two that can be realized via a Sarkisov link starting from $\mathbf{P}(1,1,4)$ or as a hypersurface of degree $5$ in $\mathbf{P}(1,1,1,4)$ (see [Nak07, Proposition 7.4, 7.11]).
@@
Let $X$ be a K3 surface admitting an involution of strictly elliptic type. Suppose that $C_{g}$ is the smooth irreducible curve of genus $2\leq g\leq 10$ fixed by $\iota$.
@@
First, let us consider $\mathcal{E}:X\to\mathbf{P}^{1}$ an arbitrary elliptic fibration. Since $\iota$ is strictly elliptic, we know that $\iota$ must map each fiber of $\mathcal{E}$ to itself (see Remark 3.3), and thus $\mathcal{E}$ factors through a fibration $f:X/\langle\iota\rangle\simeq Z_{d}\to\mathbf{P}^{1}$. Since $Z_{d}$ is a smooth del Pezzo surface and $f$ has connected fibers, it follows that $f$ is a $K_{Z_{d}}$-negative contraction with one-dimensional fibers and thus is a conic bundle by [And85, Theorem 3.1 (ii)].
@@
${X}$${\mathbf{P}^{1}}$$\scriptstyle{\mathcal{E}}$$\scriptstyle{\pi}$$\scriptstyle{\exists f}$
@@ 4. Examples and Néron-Severi lattices
$w^{2}=(a^{4}-a^{2}b^{2}+b^{4})s^{4}+\tfrac{1}{2}(-a^{3}b+a^{2}b^{2})s^{3}+\tfrac{1}{2}(a^{3}b+3a^{2}b^{2}+ab^{3}-4b^{4})s^{2}+({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}-}a^{2}b^{2}+b^{4}).$
@@ 5. Singular fibers and bisections
$I_{2}\quad[1,1],[1,1]$ $\pi$ $I_{3}\quad[2],[1,1]$$\pi$ $I_{4}\quad[2],[2]$$\pi$ $III\quad[1,{\color[rgb]{0.25390625,0.45703125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.45703125,0.01953125}\mathbf{1}}],[{\color[rgb]{0.2890625,0.56640625,0.88671875}\definecolor[named]{pgfstrokecolor}{rgb}{0.2890625,0.56640625,0.88671875}\mathbf{1}},1]$$\pi$ $IV\quad[{\color[rgb]{0.2890625,0.56640625,0.88671875}\definecolor[named]{pgfstrokecolor}{rgb}{0.2890625,0.56640625,0.88671875}\mathbf{2}}],[{\color[rgb]{0.25390625,0.45703125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.45703125,0.01953125}\mathbf{1}},1]$$\pi$
@@
Remark 5.3.
The proof of the previous result, along with Table 1, not only allows for the determination of admissible singular fibers in the induced elliptic fibration but also the complete configuration of these fibers.